OpenStudy (anonymous):

for f(x)=2x+1 and g(x)=x^2-7 find (f-g)(x) Answer's x^2 - 7/2x + 1, x = - 1/2 (the = has a line through it) x^2 -7/2x +1 2x +1/x^2 -7 2x+1/x^2-7, x = + square root 7 (the = also has a line through it and the + has a line under it. Someone can you please help! it would be a great help.

OpenStudy (triciaal):

for f(x)=2x+1 and g(x)=x^2-7 find (f-g)(x) (f-g)(x) means f(x) - g(x) 2x + 1 - (x^2 -7) -x^2 +2x +1 + 7 from the choices I think you meant f/g can you verify?

OpenStudy (anonymous):

yes f/g the f-g was on accident.

OpenStudy (anonymous):

Im confused on how to figure it out though. This stuff just confuses me.

OpenStudy (anonymous):

the 2x +1 - (x^2 -7) Would you add the 2x +1 to make it 3x - (x^2 -7)? @triciaal

OpenStudy (triciaal):

sorry having computer issues no f/g means f divided by g think about 12/3 but instead of numbers you have an expression

OpenStudy (triciaal):

another example 35/5 = 7

OpenStudy (anonymous):

No problem, so divide 2x+1? Like that?

OpenStudy (anonymous):

I'm just lost on this stuff.

OpenStudy (anonymous):

wait ok do you mean like... 2x+1/ x^2 -7?

OpenStudy (triciaal):

f(x) = 2x + 1 and g(x) = yes option c

OpenStudy (triciaal):

hint you cannot find a value for x unless you have an equation

OpenStudy (anonymous):

Ok so would all the f/g (x) stuff be like that then? Or does it all vary?

OpenStudy (triciaal):

whatever it says keep the order and do the required operation

OpenStudy (anonymous):

ok Thank you!!

OpenStudy (triciaal):

you are very welcome